# Quantum Pufferfish Privacy: A Flexible Privacy Framework for Quantum Systems

### v1

@@ V-A Properties of QPP Mechanisms
Modern guidelines for privacy frameworks [43] render properties such as convexity and post-processing (also known as transformation invariance) as basic requirements for privacy frameworks. Composability is another important property, which implies that a combination of privacy mechanisms is itself private. These properties are known to hold for the classical mutual information PP framework, and all of them, except for composability, hold for the classical PP framework; cf. [7, Theorem 2] and [3, Theorem 5.1], respectively.
@@
$\mathcal{Q}^{(k)}\coloneqq\left\{(\mathcal{R}^{(k)},\mathcal{T}^{(k)}):\begin{aligned} &\mathcal{R}^{(k)}\coloneqq(\mathcal{R}_{1},\ldots,\mathcal{R}_{k})\\ &\mathcal{T}^{(k)}\coloneqq(\mathcal{T}_{1},\ldots,\mathcal{T}_{k})\\ &\forall i\in\{1,\ldots,k\}\\ &(\mathcal{R}_{i},\mathcal{T}_{i})\in\mathcal{Q}\end{aligned}\right\}.$ (104)
@@
See Appendix C. ∎
The latter two properties above change if one considers a measurement class that comprises all possible measurements $\mathcal{M}^{k}$, as opposed to only product measurements. This is one of the main distinctions between the semi-classical and quantum cases, where, for the latter, joint measurements can infer more information and thus privacy degrades. The following theorem accounts for this latter scenario.
@@
In [10, Corollary III.3], the parallel composition of two mechanisms that satisfy $(\varepsilon_{i},\delta_{i})$-QDP for $i\in\{1,2\}$ is shown to be $(\varepsilon_{1}+\varepsilon_{2},\delta)$-QDP, where $\delta$ is given in (108). The proof technique is, however, different from ours. Property 3 of Theorem 1 also reveals that if one considers a restricted class of measurements (e.g., product measurements), then it is possible to achieve tighter privacy guarantees (namely, $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-QDP) than those obtained when allowing all joint measurements on the two systems. Also note that $\delta^{\prime}$ in (107) is independent of $\varepsilon_{1}$ and $\varepsilon_{2}$, whereas $\delta$ in (108) depends on them. Depending on the particular values that the parameters $\varepsilon_{i}$ and $\delta_{i}$ take, for $i\in\{1,2\}$, one of these aforementioned results provides sharper privacy guarantees.
@@ V-A1 Adaptive Composability
Adaptive composition refers to the case when each subsequently composed mechanism is chosen based on the outputs of the preceding ones. The goal is to quantify the overall privacy leakage at the output of the adaptively composed mechanism. This ideas has been studied in detail for classical privacy settings [2], and here we explore it for QPP.
@@
V-A2 Composability with Correlated States
In Property 3 of Theorem 1 and Proposition 3, we considered the case in which two mechanisms, composed either in parallel or adaptively, receive independent inputs (i.e., the input being $\rho^{X_{1}}\otimes\rho^{X_{2}}$ where $X_{i}\sim P_{X}\in\Theta$ for $i=\{1,2\}$, which are chosen independently). We now focus on the setting in which the inputs are correlated. The input is chosen as a separable state of the form
where $q$ represents a probability distribution with $q(z)\geq 0$ and $\sum_{z\in\mathcal{Z}}q(z)=1$, and $\omega^{z}$ and $\tau^{z}$ are quantum states for all $z\in\mathcal{Z}$ [^3] Note that (121) covers the case of having input states of the form $\sigma_{I}\coloneqq\sum_{(x,y)\in\mathcal{X}\times\mathcal{Y}}q(x,y)\ \omega^{x}\otimes\tau^{y}$ where for all $x\in\mathcal{X}$ and $y\in\mathcal{Y}$ $\omega^{x}$ and $\tau^{y}$ are states, by considering $z$ to be an index for multiple variables, i.e., setting $z=(x,y)$. One special case of interest is as follows:
$\sigma^{1}_{I}\coloneqq\sum_{z\in\mathcal{Z}}q(z)\,\omega_{1}^{z}\otimes\tau_{1}^{z}\ \textnormal{and}\ \sigma^{2}_{I}\coloneqq\sum_{z\in\mathcal{Z}}q(z)\,\omega_{2}^{z}\otimes\tau_{2}^{z},$ (123)
Corollary 1 (Composability of flexible QDP).
  1. 1.
  2. 2.
Proof:
Remark 8 (Extensions beyond flexible QDP).
$\Theta\subseteq\left\{P_{X}\in\mathcal{P}(\mathcal{X}):\begin{array}[]{cc}\forall\ (\mathcal{R},\mathcal{T})\in\mathcal{Q},\ \exists\ x,x^{\prime}\in\mathcal{X}&\\ \textnormal{s.t.}\ q_{\mathcal{R}}(x)=q_{\mathcal{T}}(x^{\prime})=1&\end{array}\right\},$ (126)
@@ V-B1 Depolarization Mechanism
$\leq\frac{(1-p)\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}{\frac{p}{d}\mathrm{Tr}\!\left[\mathrm{M}\right]}$ (133)
@@
$\varepsilon\geq\ln\!\left(1+\frac{d(1-p)\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]}{p\mathrm{Tr}\!\left[\mathrm{M}\right]}\right),$ (134)
@@
$k^{\prime}$ $\coloneqq\sup_{(\rho^{x_{1}},\rho^{x_{2}})\in\mathcal{W}_{\Theta}}\left\|\mathcal{E}(\rho^{x_{1}})-\mathcal{E}(\rho^{x_{2}})\right\|_{1}$ (143)
@@
Note that the domain knowledge encoded into the QDP framework may guide towards an improved accuracy/utility, as opposed to considering all neighboring states as secret pairs and all possible measurements. For the QDP framework without domain knowledge, [8, Theorem 3] shows that
@@
a QDP framework that has the capability to incorporate domain knowledge may cause less perturbation to the useful channel output of $\mathcal{E}$ in some cases. The rightmost inequality holds because $\mathcal{W}_{\Theta}$ includes only the neighboring pairs of states such that their occurrence has a positive probability, while $\rho\sim\sigma$ denotes all possible neighboring pairs.
@@
with the choice of the identity channel instead of $\mathcal{E}$ in Fig. 4. This occurs because $\left\|\rho-\sigma\right\|_{1}\leq 2$, with equality for pairs of orthogonal states, and $\left\|\mathrm{M}\right\|_{\infty}\leq\mathrm{Tr}\!\left[\mathrm{M}\right]$, with equality holding when considering all possible measurements. This is analogous to a version of the randomized response technique used to achieve classical local DP [48]. For a finite alphabet $\mathcal{X}$ with cardinality $|\mathcal{X}|$, the randomized response mechanism outputs the true value with probability $(1-q)$, and outputs a randomly chosen realization with probability $q/|\mathcal{X}|$. Then, if
@@
Fix $p\in[0,1]$, and the privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ with =$\mathcal{M}=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$. Let $\mathcal{E}$ be a quantum channel. Then $\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\cdot)\right)$is $\varepsilon$-QPP if
$p\geq\min\!\left\{0,\frac{d(K^{\prime}-\delta)}{dK^{\prime}+e^{\varepsilon}-1}\right\},$ (149)
@@
The proof follows from the use of the equivalent formulation through the hockey-stick divergence and the properties of this divergence. By [10, Lemma IV.I], we have
@@
We also have the property [10, Lemma II.4]
@@ V-B2 Classical PP Mechanisms from QPP Mechanisms
The QPP formalism provides a direct methodology to design classical PP mechanisms with the assistance of QPP mechanisms. In this case, we use quantum encoding to convert classical data to quantum data. We denote the quantum encoding of classical data $x\in\mathcal{X}^{n\times k}$ as $\rho^{x}\coloneqq|x\rangle\!\langle x|$. Then, we ensure the privacy for the quantum data (quantum encoding) such that the privacy is ensured for the underlying classical data.
@@
Depolarization is a common kind of noise considered in quantum information processing. Thus, this provides a method for designing classical PP mechanisms by combining the results presented in Proposition 5 and Theorem 3. It is important to note that quantum encoding of classical data would require computational resources, shifting the complexity of the mechanism design phase to the encoding phase.
@@ VI Quantifying Privacy-Utility Tradeoff
It is important to identify how the performance or utility degrades with the application of a privacy mechanism. Although the application of a privacy mechanism has the intended effect of increasing the privacy of data, it can have the detrimental effect of corrupting the data as well, thus reducing the utility of the data. Thus, it is a goal to preserve utility as much as possible while maintaining privacy. However, there is a tradeoff when trying to attain both goals. Here we define a utility metric to help quantify this tradeoff, and we prove that it can be characterized by an SDP. We then use that metric to analyse privacy-utility tradeoffs related to privacy mechanisms.
VI-A Defining Utility Metrics
We now propose a utility metric based on an operational approach. Let $\mathcal{A}$ denote a privacy mechanism. We focus on the capability of reversing the action of $\mathcal{A}$ by post-processing it with $\mathcal{B}$, in order to recover the original state input to $\mathcal{A}$ up to an error $\gamma$. Based on that operational perspective, we define $\gamma$-utility in terms of how distinguishable $\mathcal{B}\circ\mathcal{A}$ is from the identity channel, and we use the normalized diamond distance as the distinguishability measure.
@@ VI-A Defining Utility Metrics
$\mathsf{U}(\mathcal{A})\coloneqq\inf_{\mathcal{B}}\frac{1}{2}\left\|\mathcal{I}-\mathcal{B}\circ\mathcal{A}\right\|_{\diamond}\leq\gamma,$ (165)
@@
The defined utility metric can be reformulated as an SDP by using the dual SDP form of the diamond distance [49, Section 4] and rewriting the quantum channel $\mathcal{B}$ in terms of its Choi matrix, as well as translating the conditions for $\mathcal{B}$ to be a channel to conditions on its Choi matrix. In the following proposition, we employ the following facts. The Choi matrix of the composed channel $\mathcal{B}\circ\mathcal{A}$ is given by [27, Eq. (3.2.22)]
@@
and $\Gamma$ represents the Choi matrix with the subscripts showing the input and the output system of the channel while the superscript indicating the channel considered, with no superscript for the identity channel. Recall that $\mathrm{T}_{C}$ is the partial transpose operation acting on system $C$, and $\mathcal{B}$ is a quantum channel if and only if its Choi matrix $\Gamma_{CD}^{\mathcal{B}}$ satisfies $\Gamma_{CD}^{\mathcal{B}}\geq 0$ and $\mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{c}$.
Proposition 6 (SDP formulation of $\gamma$-utility).
$\mathsf{U}(\mathcal{A})=\inf_{\begin{subarray}{c}\mu\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\mu:\\ Z_{AD}\geq\Gamma_{AD}-\Gamma_{AD}^{\mathcal{B}\circ\mathcal{A}},\\ \mu I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{c}\end{array}\right\}\leq\gamma.$ (167)
The above presented SDP formulation can be used to compute the utility of a privacy mechanism $\mathcal{A}$ of interest.

### v2

@@ V-A Properties of QPP Mechanisms
Modern guidelines for privacy frameworks [54] render properties such as convexity and post-processing (also known as transformation invariance) as basic requirements for privacy frameworks. Composability is another important property, which implies that a combination of privacy mechanisms is itself private. These properties are known to hold for the classical mutual information PP framework, and all of them, except for composability, hold for the classical PP framework; cf. [12, Theorem 2] and [3, Theorem 5.1], respectively.
@@
$\mathcal{Q}^{(k)}\coloneqq$
$\left\{(\mathcal{R}^{(k)},\mathcal{T}^{(k)}):\begin{aligned} &\mathcal{R}^{(k)}\coloneqq(\mathcal{R}_{1},\ldots,\mathcal{R}_{k}),\\ &\mathcal{T}^{(k)}\coloneqq(\mathcal{T}_{1},\ldots,\mathcal{T}_{k})\\ &\forall i\in\{1,\ldots,k\}\ (\mathcal{R}_{i},\mathcal{T}_{i})\in\mathcal{Q}\end{aligned}\right\}.$ (96)
@@
See Appendix D. ∎
More broadly, parallel composition (i.e., Property 3 of Theorem 1) holds under particular separable measurements that are defined as follows:
$\left\{\sum_{j}\mathrm{M}_{1}^{(j)}\otimes\cdots\otimes\mathrm{M}_{k}^{(j)}:\forall i\sum_{j}\mathrm{M}_{i}^{(j)}\in\mathcal{M}_{i},\ \forall i,j\ \mathrm{M}_{i}^{(j)}\geq 0\right\}$ (98)
where product measurements considered in Theorem 1 are a special case.
The latter two properties of Theorem 1 change if one considers a measurement class that comprises all possible measurements ¯$\bar{\mathcal{M}}^{k}$, as opposed to only product measurements. This is one of the main distinctions between the semi-classical and quantum cases, where, for the latter, joint measurements can infer more information and thus privacy degrades. The following theorem accounts for this latter scenario.
@@
In [15, Corollary III.3], the parallel composition of two mechanisms that satisfy $(\varepsilon_{i},\delta_{i})$-QDP for $i\in\{1,2\}$ is shown to be $(\varepsilon_{1}+\varepsilon_{2},\delta)$-QDP, where $\delta$ is given in (101). The proof technique is, however, different from ours. Property 3 of Theorem 1 also reveals that if one considers a restricted class of measurements (e.g., product measurements), then it is possible to achieve tighter privacy guarantees (namely, $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-QDP) than those obtained when allowing all joint measurements on the two systems. Also note that $\delta^{\prime}$ in (100) is independent of $\varepsilon_{1}$ and $\varepsilon_{2}$, whereas $\delta$ in (101) depends on them. Depending on the particular values that the parameters $\varepsilon_{i}$ and $\delta_{i}$ take, for $i\in\{1,2\}$, one of these aforementioned results provides sharper privacy guarantees.
@@ V-A1 Adaptive Composability
Adaptive composition refers to the case when each subsequently composed mechanism is chosen based on the outputs of the preceding ones. The goal is to quantify the overall privacy leakage at the output of the adaptively composed mechanism. This idea has been studied in detail for classical privacy settings [2], and here we explore it for QPP.
@@
Remark 10 (Composability with correlated states).
In Property 3 of Theorem 1 and Proposition 3, we considered the case in which two mechanisms composed in parallel, receive independent inputs (i.e., the input being $\rho^{X_{1}}\otimes\rho^{X_{2}}$ where $X_{i}\sim P_{X}\in\Theta$ for $i=\{1,2\}$, which are chosen independently). In Appendix F we study the setting in which the inputs are correlated. There, we observe that QPP shares similar properties related to composability of classical PP frameworks, where the class of $\Theta$ plays a key role in composability to hold in general.
In Proposition 3, we assume a local structure of measurements conducted in the process, as shown in Fig. 3. This assumption is mainly motivated by technical considerations, as we can treat the resulting setting using our existing set of tools. Exploration of advanced adaptive composition techniques, which holds for more general classes of measurements is an interesting avenue for future work. In Section IX-B, we present a variant of QPP where adaptive composition holds for general measurements and strategies (refer to Fig. 8 and Remark 25).
@@ V-B1 Depolarization Mechanism
$\leq\frac{(1-p){\left|\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]\right|}}{\frac{p}{d}\mathrm{Tr}\!\left[\mathrm{M}\right]}$ (120)
@@
$\varepsilon\geq\ln\!\left(1+\frac{d(1-p){\left|\mathrm{Tr}\!\left[\mathrm{M}\!\left(\mathcal{E}(\rho^{\mathcal{R}})-\mathcal{E}(\rho^{\mathcal{T}})\right)\right]\right|}}{p\mathrm{Tr}\!\left[\mathrm{M}\right]}\right),$ (121)
@@
Note that the parameter $K$ derived from Theorem 3 represents the domain knowledge accessible and incorporated into the privacy model of the $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ QPP framework.
@@
$k^{\prime}$ $\coloneqq\sup_{(\rho^{x_{1}},\rho^{x_{2}})\in\mathcal{W}_{\Theta}}\left\|\mathcal{E}(\rho^{x_{1}})-\mathcal{E}(\rho^{x_{2}})\right\|_{1},$ (130)
@@
Note that the domain knowledge encoded into the QDP framework may guide towards an improved accuracy/utility, as opposed to considering all neighboring states as secret pairs and all possible measurements. For the QDP framework without domain knowledge, [13, Theorem 3] shows that
@@
a QDP framework that has the capability to incorporate domain knowledge may cause less perturbation to the useful channel output of $\mathcal{E}$ in some cases. The rightmost inequality holds because $\mathcal{W}_{\Theta}$ includes only the neighboring pairs of states such that their occurrence has a positive probability, while $\rho\sim\sigma$ denotes all possible neighboring pairs. Furthermore, we always have that $\frac{\left\|\mathrm{M}\right\|_{\infty}}{\mathrm{Tr}\!\left[\mathrm{M}\right]}\leq 1$ for every measurement operator $\mathrm{M}$.
@@
with the choice of the identity channel instead of $\mathcal{E}$ in Fig. 4. This occurs because $\left\|\rho-\sigma\right\|_{1}\leq 2$, with equality for pairs of orthogonal states, and $\left\|\mathrm{M}\right\|_{\infty}\leq\mathrm{Tr}\!\left[\mathrm{M}\right]$, with equality holding whenever $\mathrm{M}$ is a rank-one measurement operator. This is analogous to a version of the randomized response technique used to achieve classical local DP [59, 60, 61]. For a finite alphabet $\mathcal{X}$ with cardinality $|\mathcal{X}|$, the randomized response mechanism outputs the true value with probability $1-q$, and it outputs a randomly chosen realization with probability $q/|\mathcal{X}|$. Then, if
@@
Fix $p\in[0,1]$ and the privacy framework $(\mathcal{S},\mathcal{Q},\Theta,\mathcal{M})$ with =$\mathcal{M}=\{\mathrm{M}:0\leq\mathrm{M}\leq\mathrm{I}\}$. Let $\mathcal{E}$ be a quantum channel. Then $\mathcal{A}^{p}_{\mathrm{Dep}}\!\left(\mathcal{E}(\cdot)\right)$ is $\varepsilon$-QPP if
$p\geq\max\!\left\{0,\frac{d(K^{\prime}-\delta)}{dK^{\prime}+e^{\varepsilon}-1}\right\},$ (136)
@@
The proof follows from the use of the equivalent formulation through the hockey-stick divergence and the properties of this divergence. By [15, Lemma IV.I], we have
@@
We also have the property [15, Lemma II.4]
@@ V-B2 Classical PP Mechanisms from QPP Mechanisms
The QPP formalism provides a direct methodology to design classical PP mechanisms with the assistance of QPP mechanisms.  [^5] In the context of classical PP mechanisms, the main attempt has been the introduction of Wasserstein mechanisms, based on the infinity order Wasserstein distance and its modifications. In particular, [4] and [62] introduced these mechanisms to achieve $(\varepsilon,0)$-PP and $(\varepsilon,\delta)$-PP, respectively. However, it is important to note that these approaches may encounter computational intractability challenges. In this case, we use quantum encoding to convert classical data to quantum data. We denote the quantum encoding of classical data $x\in\mathcal{X}^{n\times k}$ as $\rho^{x}\coloneqq|x\rangle\!\langle x|$ (recall that $p_{X}\in\Theta_{c}$ are discrete probability distributions over the probability space $\mathcal{P}(\mathcal{X}^{n\times k})$ and this leads to a finite collection of $\{\rho^{x}\}_{x}$ quantum encodings). Then, we ensure the privacy for the quantum data (quantum encoding) such that the privacy is ensured for the underlying classical data.
@@
Depolarization is a common kind of noise considered in quantum information processing. Thus, this offers a method for designing classical PP mechanisms by combining the results presented in Proposition 5 and Theorem 3. However, it is essential to recognize that quantum encoding of classical data would require additional computational resources, shifting the complexity of the mechanism design phase to the encoding phase.
@@ VI Quantifying Privacy-Utility Tradeoff
In this section, we aim to assess the utility achievable through the implementation of a privatization mechanism while adhering to privacy constraints and characterize the inherent tradeoffs involved in this process. To achieve this, we define a utility metric grounded in an operational approach and demonstrate its representation via an SDP. Subsequently, we leverage this metric to conduct an in-depth analysis of privacy-utility tradeoffs, with a specific emphasis on the depolarization mechanism.
VI-A Utility Metric
Let $\mathcal{A}$ denote a privacy mechanism. We focus on assessing the potential of reversing the effects of $\mathcal{A}$ by applying a post-processing mechanism $\mathcal{B}$ to recover the initial input state to $\mathcal{A}$ up to an error $1-\gamma$. We define $\gamma$-utility in terms of how distinguishable $\mathcal{B}\circ\mathcal{A}$ is from the identity channel, employing the normalized diamond distance as the distinguishability measure.
@@ VI-A Utility Metric
$\mathsf{U}(\mathcal{A})\coloneqq 1-\inf_{\mathcal{B}}\frac{1}{2}\left\|\mathcal{I}-\mathcal{B}\circ\mathcal{A}\right\|_{\diamond}\geq\gamma,$ (152)
@@
The defined utility metric can be reformulated as an SDP by using the dual SDP form of the diamond distance [63, Section 4] and rewriting the quantum channel $\mathcal{B}$ in terms of its Choi matrix $\Gamma_{CD}^{\mathcal{B}}$, as well as translating the conditions for $\mathcal{B}$ to be a channel to conditions on its Choi matrix, namely $\Gamma_{CD}^{\mathcal{B}}\geq 0$ and $\mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{c}$.
Proposition 6 (SDP formulation of $\gamma$-utility).
The $\gamma$-utility of a privacy mechanism $\mathcal{A}$ can be formulated as the following SDP:
$\mathsf{U}(\mathcal{A})=1-\inf_{\begin{subarray}{c}\mu\geq 0\\ Z_{AD}\geq 0\\ \Gamma_{CD}^{\mathcal{B}}\geq 0\end{subarray}}\left\{\begin{array}[c]{c}\mu:\\ Z_{AD}\geq\Gamma_{AD}-\Gamma_{AD}^{\mathcal{B}\circ\mathcal{A}},\\ \mu I_{A}\geq\mathrm{Tr}_{D}\!\left[Z_{AD}\right],\\ \mathrm{Tr}_{D}\!\left[\Gamma_{CD}^{\mathcal{B}}\right]=I_{C}\end{array}\right\}\geq\gamma,$ (153)
where
@@
and $\Gamma$ represents the Choi matrix with the subscripts showing the input and the output system of the channel while the superscript indicating the channel considered, with no superscript for the identity channel. [^6] The Choi matrix of the composed channel $\mathcal{B}\circ\mathcal{A}$ is denoted by $\Gamma_{AD}^{\mathcal{B}\circ\mathcal{A}}$ and (154) follows from [37, Eq. (3.2.22)].
Note that a similar SDP formulation for approximate degradability where the identity channel in Definition 5 is replaced by the complementary channel is presented in [64, Proposition 9].
Remark 12 (Characterizing optimal privacy-utility tradeoffs).
The optimal utility attained by an $(\varepsilon,\delta)$-QPP mechanism can be characterized as an SDP. To achieve this, we combine the equivalent formulation of QPP via the DL divergence from Proposition 1 with the SDP formulation of DL divergence in Lemma 1. Combining this with the SDP formulated from Proposition 6 enables computing the privacy requirements and quantifying utility together. Additionally, we determine the optimal privacy parameters for fixed utility requirements. For a comprehensive discussion of this point, please refer to Appendix G. The utilization of the SDP derived for the DL divergence in this operational task highlights an advantage of the equivalent formulation for QPP using the DL divergence.